### v1

# DINO-Foresight: Looking into the Future with DINO

Abstract
Predicting future dynamics is crucial for applications like autonomous driving and robotics, where understanding the environment is key. Existing pixel-level methods are computationally expensive and often focus on irrelevant details. To address these challenges, we introduce DINO-Foresight, a novel framework that operates in the semantic feature space of pretrained Vision Foundation Models (VFMs). Our approach trains a masked feature transformer in a self-supervised manner to predict the evolution of VFM features over time. By forecasting these features, we can apply off-the-shelf, task-specific heads for various scene understanding tasks. In this framework, VFM features are treated as a latent space, to which different heads attach to perform specific tasks for future-frame analysis. Extensive experiments show that our framework outperforms existing methods, demonstrating its robustness and scalability. Additionally, we highlight how intermediate transformer representations in DINO-Foresight improve downstream task performance, offering a promising path for the self-supervised enhancement of VFM features. We provide the implementation code at https://github.com/Sta8is/DINO-Foresight.
1 Introduction
Predicting future states in video sequences is a key challenge in computer vision and machine learning, with important applications in autonomous systems such as self-driving cars and robotics [26, 22], which must navigate dynamic environments safely. While advances have been made in video understanding [44, 56] and generation [38], predicting future states remains difficult, especially in complex scenarios involving interactions between multiple objects over long time horizons.
Most existing approaches focus on future prediction at the pixel level, aiming to model both low-level appearance and semantic changes. These methods include generative approaches like diffusion-based models [99, 27, 38, 39, 34, 8], autoregressive models [42, 40, 52, 86], and masked video generation [95, 96]. However, for decision-making systems like self-driving cars, the primary goal is not to generate realistic-looking RGB future frames but to perceive a scene and predict its evolution over time at a more semantic or abstract level, such as identifying what objects exist in the scene and where. Thus, future prediction approaches defined at the pixel space may allocate excessive capacity to modeling low-level appearance, which may be irrelevant to the scene understanding needed for decision-making.
Previous work has shown that directly predicting task-specific features for future frames can be more effective than generating full RGB frames and then performing tasks like segmentation or depth estimation [62]. However, existing methods have limitations: some are designed for a single task (e.g., segmentation) [66, 18, 79], making it difficult to handle multiple tasks simultaneously, while others attempt to predict features for multiple tasks at once [41], resulting in complex architectures that do not scale well.
To address these challenges, we leverage recent advances demonstrating the effectiveness of features derived from large-scale Vision Foundation Models (VFMs) for diverse scene understanding tasks, with strong generalization across varied [9, 68, 69]. Building on this foundation, in this work we propose a novel approach to future prediction that operates directly within the space of VFM features. Specifically, we introduce a masked feature transformer architecture, which processes VFM features from previous frames to predict their future counterparts. Using these features, our method inherits a strong and versatile understanding of the scene, offering a unified semantic representation of future states. This representation can be readily adapted to various tasks through task-specific prediction heads applied to the predicted features (see Fig. 1).
Our method offers several key advantages:
Simplicity, scalability, and semantic focus:
It leverages the benefits of generative pixel-level methods, providing simplicity, self-supervised training, and scalability, while avoiding the need to model low-level appearance details.
Task-agnostic:
Unlike prior task-specific semantic future prediction methods, our approach is flexible and can be applied to any scene understanding task, without retraining the model when task requirements change.
Plug-and-play modular task heads:
We use a modular approach where different task-specific heads can be easily added or removed from the framework without retraining the future feature prediction model.
Our contributions are fourfold: (1) We introduce DINO-Foresight, a self-supervised method for semantic future prediction that leverages VFM features. This approach is scalable and can be applied to a variety of scene understanding tasks in a straightforward and practical way. (2) We develop an encoder-decoder transformer architecture for VFM feature forecasting, using a pretrained VFM as the encoder and a masked feature transformer as the decoder that performs the next-frame feature prediction task (see Fig. 1). This setup efficiently propagates multi-layer and high-resolution features, which is crucial for scene understanding task performance. (3) We demonstrate the effectiveness of our approach through extensive experiments, showing superior performance in multiple scene understanding tasks, including semantic and instance segmentation, depth estimation, and surface normals prediction. (4) Last, but not least, we show that the intermediate features within our masked transformer can further improve the performance of downstream tasks. This finding illustrates the promise of our method as a self-supervised visual learning strategy that enhances the already strong VFM features.
2 Related Work
Figure 1: Future Prediction with Masked Feature Transformer. Our framework trains a masked transformer to predict the temporal evolution of features extracted by a frozen visual encoder based on context frames in a self-supervised manner, minimizing SmoothL1 loss between predicted and future features. By forecasting these rich and versatile features, task-specific prediction heads—such as semantic segmentation, depth, and surface normals—can be effortlessly employed at test time, enabling modular and efficient scene understanding.
Future Prediction  Video future prediction and generation hold significant promise for applications like autonomous driving and robotics, where crafting future frames from past observations is crucial yet challenging due to the high dimensionality and unpredictability of video data. Traditional methods like Convolutional Long Short-Term Memory networks (Conv-LSTMs) [66, 91, 87, 10, 54, 89, 28] have grappled with maintaining visual realism and consistency over time. Subsequent advancements have leveraged generative adversarial networks (GANs) and variational autoencoders (VAEs) [92, 84, 10, 53, 3], alongside diffusion models [38, 39, 27, 34], to enhance spatial-temporal coherence and improve the quality of predictions. Furthermore, transformer-based models have also been adapted to videos, utilizing auto-regressive and masked modeling objectives to capture video dynamics [96, 95, 33, 86].
Future Feature Prediction  An emerging approach for future-frame prediction focuses on forecasting intermediate features from an encoder rather than predicting raw RGB values [66, 74, 77, 15, 45, 63, 73, 43, 83, 100]. This strategy models the abstract representations learned by the encoder, which are then used by task-specific heads for downstream tasks. While effective, these methods often rely on encoders specialized for specific tasks or datasets, limiting their generalizability across domains. To address this, we use encoders from visual foundation models, which, with their large-scale pre-training, perform well across a variety of tasks and generalize effectively to new scenes.
Vision Foundation Models (VFMs)
VFMs have transformed computer vision, achieving strong performance across a range of visual tasks. Trained on large-scale datasets, these models learn rich, transferable visual representations. Notable examples include DINOv2 [9, 68], a self-supervised model based on self-distillation; CLIP and its variants [69, 24, 78, 25], which align visual representations with natural language; and SAM (Segment Anything Model) [50], a foundation model for object segmentation. In this work, we explore the potential of VFM features for semantic future prediction tasks, connecting static visual understanding with dynamic prediction.
Multi Task Learning  Multi-Task Learning (MTL) is a learning paradigm that enables simultaneous training of models on multiple related tasks [64, 65, 67, 82], promoting shared representations and improving performance across tasks. Traditional MTL frameworks often use parameter sharing [48, 75, 5] or task interaction allowing exchange of information [7, 14, 16, 65, 71]. Other approaches employ a strategy that incrementally increases the model’s depth during training, enabling the network to learn task-specific representations in a more resource-efficient way [1, 19, 32, 61, 97]. Recently, the emergence of large-scale pretrained models has led to the introduction of adapter-based multi-task fine-tuning approaches [55, 59]. In our work, we leverage VFM features to provide a unified, scalable and modular framework for future prediction. Our approach enables seamless integration of multiple tasks without retraining or complex adaptations.
3 Methodology
Our semantic future prediction framework leverages VFMs trained on large-scale data, which have proven highly effective in various scene understanding tasks and in generalizing to unseen scenes. We propose training a masked feature transformer model in a self-supervised manner to predict the evolution of these VFM features across future frames. By forecasting these features, we can apply various off-the-shelf, task-specific heads for scene understanding. In this approach, we treat the feature representations of VFMs as a latent space, to which different heads can be connected to perform specific tasks for future-frame analysis.
Fig. 1 provides an overview of our approach, with the key components detailed in the following sections: Section 3.1 describes how target features are generated from multi-layer VFM features and the strategies used to manage their high dimensionality. Section 3.2 explains how we formulate the future-frame prediction task as a masked feature modeling problem and the transformer architecture used. Section 3.3 covers techniques for efficiently training the transformer to predict high-resolution VFM features. Finally, Section 3.4 presents our modular framework for multi-task predictions using future-frame features and details how prediction heads are trained and integrated.
3.1 Hierarchical Target Feature Construction
In Fig. 2, we provide an overview of how the target feature space for the masked feature transformer is constructed. Below, we outline the main steps involved.
Multi-Layer VFM Feature Extraction  Scene understanding models often benefit from processing features across multiple layers of an image encoder [70, 60, 57, 17, 98, 12], especially when the encoder is frozen, as in our approach. To fully leverage the pretrained representations of VFMs, we propagate features extracted from multiple layers of the VFM.
Our framework uses VFMs based on the Vision Transformer (ViT) architecture, though the approach can be extended to other architectures. Given a sequence of $N$ image frames $\mathbf{X}\in\mathbb{R}^{N\times H^{\prime}\times W^{\prime}\times 3}$, let $\mathbf{F}^{(l)}\in\mathbb{R}^{N\times H\times W\times D_{enc}}$ represent the features extracted from layer $l$ of the ViT model. Here, $D_{enc}$ is the feature dimension, and $H\times W$ is the spatial resolution, which are consistent across layers in ViT-based models. To form the target feature space on which the masked feature transformer operates, we concatenate the features from $L$ layers along the channel dimension, resulting in $\mathbf{F}_{\text{concat}}\in\mathbb{R}^{N\times H\times W\times L\cdot D_{enc}}$. These concatenated features capture rich semantic information from the input images at multiple levels of abstraction.
Dimensionality Reduction
The concatenated features $\mathbf{F}_{\text{concat}}$ have high dimensionality, so we apply dimensionality reduction to simplify the prediction task for the transformer while retaining essential information. In this work, we use Principal Component Analysis (PCA) to reduce the dimensionality, transforming $\mathbf{F}_{\text{concat}}$ into a lower-dimensional representation $\mathbf{F}_{\text{PCA}}\in\mathbb{R}^{N\times H\times W\times D}$, where $D\ll L\cdot D_{enc}$. These PCA-reduced features, $\mathbf{F}_{\text{PCA}}$, serve as the target features on which the masked feature transformer operates, i.e., $\mathbf{F}_{\text{TRG}}=\mathbf{F}_{\text{PCA}}$. While PCA is used in this work, alternative methods, such as Variational Autoencoders (VAE), could also be explored in future work.
3.2 Self-Supervised Future Feature Prediction with Masked Transformers
Figure 2: Hierarchical Target Feature Construction for Masked Feature Transformer. Our framework builds a feature space by extracting and concatenating multi-layer features from a frozen ViT encoder, capturing semantic information at different levels of abstraction. To reduce their high dimensionality, we apply PCA, creating more compact features for the masked feature transformer to operate on.
Masked Feature Transformer Architecture
Inspired by previous video generation models [11, 33], we implement the future feature prediction task using a self-supervised masked transformer architecture. The task involves predicting future frames in a video sequence consisting of $N$ frames, where $N_{c}$ are context frames and $N_{p}$ are future frames to be predicted, such that $N=N_{c}+N_{p}$. Given the target features $\mathbf{F}_{\text{TRG}}\in\mathbb{R}^{N\times H\times W\times D}$ for these $N$ frames, the future-frame tokens are masked, and the transformer must predict these missing tokens by processing all the tokens from the entire sequence, i.e., all the $N\cdot H\cdot W$ tokens.
The transformer architecture begins with a token embedding stage, where each token is projected from $D$ dimensions into the transformer’s hidden dimension $D_{dec}$ through a linear layer. During training, the tokens corresponding to the future frames are replaced with a learnable $D_{\text{dec}}$-dimensional [MASK] vector. During inference, these [MASK] tokens are appended after the context frames. Each token also receives a position embedding to retain both temporal and spatial information across the sequence.
The tokens are then passed through a series of transformer layers. Standard self-attention layers in transformers have quadratic time complexity with respect to the number of tokens, making them computationally expensive for high-resolution, multi-frame sequences. To address this, we follow the approach from recent video transformers [2, 33] and decompose the attention mechanism into temporal and spatial components. Temporal attention is applied across tokens with the same spatial position in different frames, capturing the dynamics over time. Spatial attention, on the other hand, operates within individual frames, focusing on spatial interactions. Thus, each transformer layer consists of a temporal Multi-Head Self-Attention (MSA) layer, a spatial MSA layer, and a feedforward MLP layer. After passing through the transformer layers, a linear prediction layer maps the output token embeddings from the hidden dimension $D_{\text{dec}}$ back to the feature dimension $D$, producing the predicted feature map $\mathbf{\tilde{F}}_{\text{TRG}}\in\mathbb{R}^{N\times H\times W\times D}$.
Training Objective for Masked Feature Modeling
We frame the future-frame prediction as a continuous regression problem and optimize a self-supervised training objective based on the SmoothL1 loss between the predicted features $\mathbf{\tilde{F}}_{\text{TRG}}$ and the ground truth features $\mathbf{F}_{\text{TRG}}$ at the masked locations.
The loss function is defined as:
$\mathcal{L}_{\mathrm{MFM}}=\underset{x\in\mathcal{X}}{\mathbb{E}}\left[\sum_{p\in\mathcal{P}}M(p)\cdot L\left(\mathbf{F}_{\text{TRG}}(p),\mathbf{\tilde{F}}_{\text{TRG}}(p)\right)\right],$ (1)
where $\mathcal{X}$ denotes the training dataset, $\mathcal{P}=[1:N]\times[1:H]\times[1:W]$ represents the set of positions across the $N$, $H$, and $W$ dimensions, $\mathbf{F}_{\text{TRG}}(i)$ and $\mathbf{\tilde{F}}_{\text{TRG}}(i)$ are the ground truth and predicted feature vectors at position $p$, respectively, and $M$ is a binary mask that is non-zero only at the positions corresponding to future frames. The function $L(\cdot,\cdot)$ computes the SmoothL1 loss between two feature vectors:
$L(x,y)=\sum_{d=1}^{D}\begin{cases}0.5\frac{(x_{d}-y_{d})^{2}}{\beta},&\text{if}|x_{d}-y_{d}|<\beta,\\ |x_{d}-y_{d}|-0.5\beta,&\text{otherwise.}\end{cases}$ (2)
In our experiments, we set $\beta=0.1$.
3.3 Compute-Efficient Training Strategies for High-Resolution Feature Forecasting
Using high-resolution features is crucial for pixel-wise scene understanding tasks, such as segmentation or depth prediction, where low-resolution features struggle to capture small objects or fine spatial structures [70, 17, 12, 81] To achieve good performance on these tasks, we aim to forecast VFM features extracted from frames with a spatial resolution of $H^{\prime}\times W^{\prime}=448\times 896$. For a ViT with a patch size of $14\times 14$, as used in DINOv2 [68] and EVA-CLIP [25], this results in feature maps with a resolution of $H\times W=32\times 64$, corresponding to 2048 tokens per frame.
However, training ViTs on such high-resolution inputs is computationally expensive in terms of both time and memory [23]. To address this challenge while maintaining efficient training, we explore the following strategies:
Low-Resolution Training with High-Resolution Inference
In this approach, we train on frames with a lower resolution of $224\times 448$, resulting in features with a resolution of $16\times 32$. During testing, we use high-resolution frames ($448\times 896$) and adapt the position embeddings through interpolation. However, this strategy leads to suboptimal performance due to a distribution shift between the training and test data, which causes inaccurate feature forecasting.
Sliding-Window Approach for High-Resolution Inference
Inspired by sliding-window techniques used in segmentation tasks [76], this strategy trains the model with cropped feature maps. The ViT encoder extracts features from high-resolution frames ($448\times 896$), producing high-resolution tokens (e.g., $32\times 64$ for a patch size of $14\times 14$). During training, we sample local crops of size $16\times 32$, taken from the same spatial locations across frames. The masked transformer is trained on these cropped features. During inference, the model applies the transformer in a sliding window fashion to the high-resolution features using these same local crops. This approach enables the model to efficiently handle large inputs without the computational burden of training on high-resolution features.
Two-Phase Training with Resolution Adaptation
Inspired by  [68, 81, 51], this strategy employs a two-phase training process. First, the model is trained on low-resolution frames ($224\times 448$) for several epochs, focusing on learning broad feature forecasting. Then, the model is fine-tuned on high-resolution frames ($448\times 896$) for a small number of epochs. This adaptation phase improves the model’s ability to handle high-resolution features without incurring the computational cost of training from scratch at the higher resolution.
As shown in our experiments, both strategies are effective, but the two-phase approach yields better feature forecasting performance. This is likely because the masked transformer has access to a larger spatial context when propagating VFM features in future frames.
3.4 Modular Framework for Future-Frame Multi-Task Predictions
Our framework is designed to work with a library of task-specific prediction heads, which can be easily applied to predicted future features after self-supervised training of the masked feature transformer. This approach is modular with respect to task heads: each head serves a different task and can be added or removed as needed without requiring retraining of the masked feature transformer.
We focus on three pixel-wise prediction tasks: semantic segmentation, depth prediction, and surface normals estimation. However, our framework can also support other scene understanding tasks, such as object detection and instance segmentation. For these tasks, we use the Dense Prediction Transformer [70] (DPT) architecture, which is well-suited to our setup. DPT is designed to work with multi-layer features from ViT-based encoders, aligning with the type of features that our masked feature transformer model learns to predict. DPT assembles these multi-layer ViT features, processes them at different resolutions, and progressively combines them into high-resolution predictions using convolutional layers. While we focus on DPT, our framework is flexible and can also accommodate other prediction heads, such as Mask2Former [17] or Mask R-CNN [36].
Prediction heads can be trained directly on frozen VFM features and then applied “off-the-shelf” to future-frame features predicted by the masked feature transformer. Additionally, they can be trained to account for the PCA stage by applying PCA compression and decompression to the multi-layer features. This approach is useful for cases where prediction heads are trained on annotated 2D images, without requiring video data, and then added to the library for future-frame predictions. Alternatively, prediction heads can be trained directly on the future-frame features predicted by an already pre-trained masked feature transformer, allowing them to adapt to potential distribution shifts between the original VFM features and the forecasted VFM features. In practice, we observed only minor performance differences between these two approaches.
4 Experiments
4.1 Experimental setup
Data.  We assess our approach using the Cityscapes dataset [20], which offers video sequences of urban driving environments. The dataset includes 2,975 training sequences, 500 for validation, and 1,525 for testing, each with 30 frames captured at 16 fps and a resolution of 1024 × 2048 pixels. The 20th frame in each sequence is annotated for semantic segmentation with 19 classes. We train our model using the training sequences and evaluate it on the validation set.
Implementation details.  By default, we use DINOv2-Reg with ViT-B/14 as the default VFM visual encoder for our method. For the masked feature transformer we built upon [6] implementation. We use 12 layers with a hidden dimension of $d=1152$ and sequence length $N=5$ (with $N_{c}=4$ context frames and $N_{p}=1$ future frame). For end-to-end training, we use the Adam optimizer [49] with momentum parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of $1.6\times 10^{-4}$ with cosine annealing. Training is conducted on 8 A100 40Gb GPUs with an effective batch size of 64. We train DPT heads for the semantic segmentation, depth prediction, and surface normals estimation tasks, and a Mask2Former head for the instance segmentation task. More implementation details Appendix C.
Evaluation Metrics.  To evaluate our method’s performance in predicting future semantic segmentation and depth maps, we use the following metrics: For semantic segmentation, we use mean Intersection over Union (mIoU) in two ways: (1) mIoU (ALL), which includes all semantic classes, and (2) MO-mIoU (MO), which considers only movable object classes like person, rider, car, truck, bus, train, motorcycle, and bicycle. For instance segmentation, we measure performance using average precision at a 0.50 IoU threshold (AP50) and also the mean average precision over IoU thresholds from 0.50 to 0.95. For depth prediction, we use two metrics: the mean Absolute Relative Error (AbsRel), defined as $\frac{1}{M}\sum_{i=1}^{M}\frac{|a_{i}-b_{i}|}{b_{i}}$, where $a_{i}$ and $b_{i}$ are the predicted and ground truth disparities at pixel $i$, and $M$ is the number of pixels. We also evaluate depth accuracy using $\delta_{1}$, the percentage of pixels where $\max\left(\frac{a_{i}}{b_{i}},\frac{b_{i}}{a_{i}}\right)<1.25$. For surface normal evaluation, we compute the mean angular error $\downarrow$ as $\frac{1}{N}\sum_{i=1}^{N}\cos^{-1}\left(\frac{\mathbf{n}_{i}\cdot\tilde{\mathbf{n}}_{i}}{\|\mathbf{n}_{i}\|\,\|\tilde{\mathbf{n}}_{i}\|}\right)$, where $\mathbf{n}_{i}$ and $\tilde{\mathbf{n}}_{i}$ are the predicted and ground truth normals, respectively. Furthermore, we measure precision through the percentage of pixels with angular errors below 11.25°, calculated as $\left(\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\theta_{i}<11.25^{\circ})\right)$.
Evaluation Scenarios  Following previous work [62, 66], we assess our model in two scenarios: short-term prediction (3 frames ahead,  0.18s) and mid-term prediction (9 frames ahead,  0.54s). In both cases, the target is the 20th frame. We subsample sequences by a factor of 3 before input to the model. For short-term prediction, the model uses frames 8, 11, 14, and 17 as context to predict frame 20 (with context length $N_{c}=4$ and $N_{p}=1$). For mid-term prediction, the model uses frames 2, 5, 8, and 11 as context and predicts frame 20 auto-regressively through frames 14 and 17. We calculate segmentation metrics on the 20th frame using Cityscapes ground truth. For depth and surface normals, we rely on pseudo-annotations from DepthAnythingV2 [94] and Lotus [35], respectively, due to the lack of true annotations in Cityscapes.
Method Short-term Mid-term
ALL MO ALL MO
3Dconv-F2F [18] 57.0 - 40.8 -
Dil10-S2S [62] 59.4 55.3 47.8 40.8
F2F [63] - 61.2 - 41.2
DeformF2F [73] 65.5 63.8 53.6 49.9
LSTM M2M [79] 67.1 65.1 51.5 46.3
APANet [43] - 64.9 - 51.4
F2MF [74] 69.6 67.7 57.9 54.6
PFA [58] 71.1 69.2 60.3 56.7
DINO Oracle 77.0 77.4 77.0 77.4
DINO-Foresight (ours) 71.8 71.7 59.8 57.6
Table 1: Comparison results with future semantic segmentation methods on Cityscapes validation set. ALL: mIoU of all classes. MO: mIoU of movable objects.
Method Short-term Mid-term
AP50 AP AP50 AP
Mask R-CNN Oracle 65.8 37.3 65.8 37.3
Mask H2F [63] 25.5 11.8 14.2 5.1
F2F [63] 39.9 19.4 19.4 7.7
CPConvLSTM [77] 44.3 22.1 25.6 11.2
APANet [43] 46.1 23.2 29.2 12.9
PFA [58] 48.7 24.9 30.5 14.8
DINO Oracle 56.1 32.1 56.1 32.1
DINO-Foresight (ours) 44.8 23.0 26.4 11.1
Table 2: Comparison with prior work on instance segmentation forecasting. DINO-Foresight and DINO Oracle use a Mask2Former prediction head.
Method Short-term Mid-term
ALL MO ALL MO
Segmenter [76] Oracle 78.6 80.8 78.6 80.6
Segmenter Copy-Last 55.5 52.7 40.5 32.2
VISTA + Segmenter 45.7 43.8 40.4 35.6
DINO Oracle 77.0 77.4 77.0 77.4
DINO Copy-last 54.7 52.0 40.4 32.3
DINO-Foresight (ours) 71.8 71.7 59.8 57.6
Table 3: Future prediction: VFM features vs RGB pixels. Results on semantic segmentation forecastig.
Encoder Method Segmentation Depth Surface Normals
Short Mid Short Mid Short Mid
[…]
DINOv2 [68] Oracle 77.0 77.4 77.0 77.4 89.1 .108 89.1 .108 3.24 95.3 3.24 95.3
Copy Last 54.7 52.0 40.4 32.3 84.1 .154 77.8 .212 4.41 89.2 5.39 84.0
Prediction 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
EVA2-CLIP [25] Oracle 71.0 69.5 71.0 69.5 85.2 .123 85.2 .123 3.37 94.5 3.37 94.5
Copy Last 51.9 47.7 38.5 29.5 81.2 .161 75.6 .216 4.52 88.5 5.44 83.6
Prediction 66.3 64.2 54.5 49.6 85.1 .122 82.5 .145 3.56 93.4 4.18 90.1
SAM [50] Oracle 69.8 63.9 69.8 63.9 84.8 .143 84.8 .143 3.01 96.0 3.01 96.0
Copy Last 49.4 41.8 36.8 26.0 78.3 .211 73.4 .267 4.84 87.4 5.77 82.4
Prediction 65.3 59.3 52.5 43.9 81.3 .178 77.6 .209 3.80 92.8 4.49 89.2
Table 4: Comparison of VFM encoders across tasks. For each encoder (DINOv2, EVA2-CLIP, SAM), we show performance on segmentation (ALL, MO), depth estimation ($\delta_{1}$ accuracy, AbsRel error), and surface normal prediction (m, percentage within 11.25°).
Resolutions Adaptation Short-term Mid-term
[…]
Oracle
(a) 224$\rightarrow$224 N/A 68.24 66.41 68.24 66.41
(b) 448$\rightarrow$448 N/A 76.97 77.40 76.97 77.40
Forecasting
(c) 224$\rightarrow$224 N/A 64.50 62.63 55.49 52.62
(d) 224$\rightarrow$448 Pos. interp. 64.34 64.29 48.31 44.60
(e) 224$\rightarrow$448224 Sliding win. 71.26 71.11 58.75 56.78
(f) (224&\&&448$\rightarrow$448 Two-phase 71.81 71.71 59.78 57.65
Table 5: Strategies for Training-Efficient High-Resolution Feature Forecasting.
4.2 VFM feature forecasting results
Comparison with Prior Work  In Tables 1 and 2, we compare our method to previous approaches in semantic and instance segmentation forecasting. On semantic segmentation forecasting, our method achieves superior results on all but one metric. For instance-segmentation forecasting, our method achieves performance comparable to the previous state-of-the-art, despite operating under a lower upper performance bound due to a weaker oracle (Mask R-CNN Oracle for prior work vs. DINO Oracle for us). This is likely due to the challenges of learning a strong instance segmentation model with a frozen visual encoder.
Unified Representations for Multiple Tasks: VFM Features vs. RGB Pixels
Our approach enables multiple scene understanding tasks with a single feature forecasting model, eliminating the need for task-specific models in previous methods, achieving competitive or superior results in semantic and instance segmentation forecasting (see Tables 1 and 2), as well as in depth prediction and surface normal estimation (see Table 4), underscoring its flexibility and practicality.
An alternative to forecasting VFM features is to predict future frames directly in RGB space, which also supports various downstream tasks. In this case, generated RGB frames can be processed with any scene understanding model. To compare these two approaches, we use VISTA [27], a recent large-scale video diffusion model for driving scenes. Given three context frames, VISTA generates 22 future frames. Using VISTA, we synthesize short-term and mid-term future frames and apply Segmenter [76] to these frames. Despite Segmenter’s superior oracle performance over our DPT head with frozen-DINOv2 features, VISTA achieves low semantic segmentation forecasting scores—even lower than Copy-Last (see Table 3). We provide an extended evaluation of our approach against the Copy-last and Vista baselines for depth and surface normals in Appendix Tab. 7. The Vista baseline only exceeds the Copy-last approach in the depth modality, while it produces poorer results in the surface normals modality.
Visual inspection (see Appendix Figs. 5 and 6) shows that VISTA achieves impressive quality and consistency for static parts of scenes or when there is relatively simple movement (e.g., all vehicles moving forward). However, it struggles with more complex motion, such as when vehicles move in different directions. This difficulty likely stems from VISTA’s need to model both low-level image details and high-level semantic changes, for which even its large-scale training proves insufficient (it was trained on 1740 hours of driving videos over 8 days on 128 A100 GPUs). In contrast, our method, which forecasts directly in the VFM feature space, abstracts away low-level pixel variations, enabling it to better forecast the semantic structure of the scene with far fewer training resources.
Comparison of VFM Visual Encoders  In Table 4, we evaluate our method using three distinct VFM encoders to extract the features on which our masked feature transformer operates (Prediction rows): DINOv2 with registers  [21, 68] (a self-supervised method), EVA2-CLIP [25] (a vision-language contrastive method), and SAM [50](a supervised instance segmentation method). For each, we use the ViT-B model variant. We also include results for the Copy-Last and Oracle baselines for comparison.
The results show that (1) DINOv2 consistently outperforms the other encoders across all tasks, providing the best results for both short-term and mid-term predictions. (2) This is in line with expectations, as the DINOv2-based Oracle also achieves the highest performance in all cases. (3) Additionally, our masked feature transformer proves effective in predicting future-frame features for all VFM encoders, showing significant improvement over the Copy-Last baseline. Based on these findings, we select DINOv2 as the default VFM visual encoder for our method.
Predicted + Intermediate Transformer FeaturesPredicted VFM features only
6910111271.572.0Transformer LayermIoU(a) Short-term, Segm. 6910111259.560.0Transformer Layer(b) Mid-term, Segm 69101112.000.002.004Transformer LayerAbsRel Reduction(c) Short-term, Depth 69101112.000.002.004Transformer Layer(d) Mid-term, Depth
Figure 3: Impact of Intermediate Transformer Features on Future Segmentation and Depth Prediction. Results are shown for semantic segmentation and depth prediction heads using two feature sets: only the VFM features predicted by the masked feature transformer (dashed line) and combined features from both predicted and intermediate transformer layers (blue bars). We evaluate DPT heads trained on features from the 6th, 9th, 10th, 11th, and 12th layers. For segmentation (barplots (a) and (b)), we report mIoU across all classes. For depth (barplots (c) and (d)), we show the reduction in AbsRel metric (higher is better) when adding intermediate layer features.
Scene Frankfurt (1) Scene Munster (15)
Last Context Frame Future Frame Last Context Frame Future Frame
Refer to caption Refer to caption Refer to caption Refer to caption
Ours Oracle Ours Oracle
Segm. Refer to caption Refer to caption Refer to caption Refer to caption
Depth Refer to caption Refer to caption Refer to caption Refer to caption
Surf. Normals Refer to caption Refer to caption Refer to caption Refer to caption
(a) Short-Term
Last Context Frame Future Frame Last Context Frame Future Frame
Refer to caption Refer to caption Refer to caption Refer to caption
Ours Oracle Ours Oracle
Segm. Refer to caption Refer to caption Refer to caption Refer to caption
Depth Refer to caption Refer to caption Refer to caption Refer to caption
Surf. Normals Refer to caption Refer to caption Refer to caption Refer to caption
(b) Mid-Term
Figure 4: Visualization of future predictions for semantic segmentation, depth, and surface normals. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Training-Efficient Strategies for High-Resolution Feature Forecasting  In Table 5, we compare the resolution-adaptation strategies from Section 3.3, reporting results for future semantic segmentation. High-resolution features are essential: comparing the low-resolution Oracle baseline (model (a)) with the high-resolution Oracle baseline (model (b)) highlights the importance of high-resolution features for strong segmentation performance. Consequently, forecasting low-resolution features (model (c)) results in significantly poorer segmentation than models predicting high-resolution features (models (e) and (f)).
Adapting a model trained on low-resolution features for high-resolution inputs by simply adjusting position embeddings during inference (model (d)) leads to suboptimal results, even underperforming compared to low-resolution forecasting. The other two adaptation strategies—Sliding Window (model (e)) and two-phase training with resolution increase (model (f))—achieve considerably better results, demonstrating their effectiveness. The two-phase approach is simpler and yields the best performance, so we adopt it as our default for high-resolution feature forecasting.
4.3 Qualitative results.
In Figure 4, we present qualitative results from our method applied to semantic segmentation, depth estimation, and surface normal prediction tasks, with both short-term and mid-term future predictions. Our single VFM feature prediction model produces meaningful outputs across all tasks, demonstrating the benefits of leveraging the feature space of large-scale pre-trained VFMs for future prediction.
4.4 Emerging Visual Representations in the Future-Frame Masked Feature Transformer
Self-supervised representation learning has achieved remarkable progress, with numerous studies focusing on extracting robust visual features from unlabeled images and videos [4, 80, 30, 72, 85, 13, 9, 37, 31, 88, 47, 46, 29]. Inspired by these advancements, we investigate the potential of our future-frame masked feature transformer as a self-supervised method for enhancing VFM visual features. Specifically, we train DPT heads for semantic segmentation and depth prediction, using not only the features predicted by the masked feature transformer but also additional features extracted from intermediate transformer layers. We examine features from the 6th, 9th, 10th, 11th, and 12th (last) layers of the transformer to assess whether these intermediate representations can further improve the strong VFM features predicted by our masked transformer.
Results, shown in Fig. 3, indicate that intermediate features from the transformer improve performance on both segmentation and depth forecasting tasks, except in one case (6th-layer features for short-term segmentation prediction). For segmentation, the best results are achieved with features from the 9th layer, while for depth, features from the 6th layer perform best. Although these improvements are modest, this aligns with expectations given the strength of the predicted VFM features alone.
While exploring self-supervised learning was not the primary aim of our work, we find these results intriguing, as they suggest that future prediction methods hold promise as self-supervised visual representation learners. We hope this work can spark further research into this direction.
5 Conclusion
We introduced DINO-Foresight, a future prediction framework that forecasts VFM features instead of raw RGB frames. Our encoder-decoder transformer architecture efficiently propagates multi-layer, high-resolution VFM features, enabling strong performance across a range of scene understanding tasks. By leveraging these features, the masked transformer predicts future state representations for multiple tasks, including semantic segmentation, instance segmentation, depth estimation, and surface normal prediction. Extensive evaluations show our method’s excellent performance, with intermediate transformer features further enhancing results, underscoring the potential of our approach for self-supervised visual representation learning.
Future work could explore scaling our method with larger, more diverse datasets and incorporating stronger VFMs like DINOv2-Huge or multiple VFMs with varied scene understanding capabilities. Additionally, exploring action conditioning to enhance controllability in autonomous driving is another promising direction for future research.
Acknowledgements
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Hardware resources were granted with the support of GRNET. Also, this work was performed using HPC resources from GENCI-IDRIS (Grants 2023-A0141014182, 2023-AD011012884R2, and 2024-AD011012884R3).
Appendix A Additional Results
A.1 Impact of PCA
In Table 6, we evaluate the impact of applying PCA to reduce the dimensionality of the multi-layer VFM features. We see that using PCA improves the semantic segmentation forecasting results for short-term prediction and especially of moving objects. However, on mid-term prediction, the differences are eliminated. In depth forecasting, PCA enhances performance across all metrics for both short-term and mid-term predictions.
Method Segmentation Depth
Short Mid Short Mid
[…]
Without PCA 71.3 70.4 59.9 57.6 87.9 .122 84.8 .147
With PCA 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136
Table 6: Impact of PCA. Results on semantic segmentation and depth forecasting.
Method Segmentation Depth Surface Normals
Short Mid Short Mid Short Mid
[…]
Oracle 77.0 77.4 77.0 77.4 89.1 .108 89.1 .108 3.24 95.3 3.24 95.3
Copy Last 54.7 52.0 40.4 32.3 84.1 .154 77.8 .212 4.41 89.2 5.39 84.0
VISTA [27] 45.7 43.8 40.4 35.6 84.9 .143 80.2 .183 4.87 85.9 5.63 83.2
DINO-Foresight 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
Table 7: Comparison across tasks. We use DINOv2 encoder and show performance on segmentation (ALL, MO), depth estimation ($\delta_{1}$ accuracy, AbsRel error), and surface normal prediction (m, percentage within 11.25°). For the VISTA baseline, we used DepthAnythingV2 [94] and Lotus [35] to generate depth and surface normal predictions respectively.
A.2 More Visualizations
In Figs. 5 and 6, we present additional qualitative results illustrating the prediction of semantic segmentation, depth maps, and surface normals. Specifically, we compare our method, DINO-Foresight, against the Oracle, which involves using future RGB frames as inputs for different prediction heads, as well as Vista [27]. For Vista, the outputs are derived from the Segmenter [76], DepthAnythingV2 [94], and Lotus [35], applied to RGB frames predicted by Vista. In these two scenes, Vista performs well in capturing static elements when the vehicle is moving forward, as illustrated in Fig. 5. However, its quality diminishes when dealing with moving objects and smaller structures, as shown in Fig. 6. In both cases, DINO-Foresight excels at maintaining the integrity of motion dynamics and accurately capturing intricate details.
In Figs. 7 and 6, we offer additional qualitative results derived from utilizing DINO-Foresight for the prediction of semantic segmentation and depth maps and surface normals over extended time intervals. These outcomes are achieved through the use of autoregressive rollouts. Beginning with a series of four context frames ($X_{t-9}$ to $X_{t}$), the model is capable of predicting up to 48 subsequent frames, equivalent to 2.88 seconds, with predictions occurring at an interval of every third frame. Our model consistently delivers accurate predictions over the entire forecasted duration, effectively capturing motion dynamics and maintaining consistency across different modalities. This performance underscores its robustness and versatility, which are related to its capability of predicting the features of a foundation model. As a final remark, it is important to note that the noisy segmentation predictions observed at the bottom of the images, present in both the predicted and Oracle results, are attributed to unannotated regions in the Cityscapes dataset that are disregarded during DPT training. This artifact impacts only the segmentation outcomes of DPT head and does not affect the predicted future features, as evidenced by the clear and accurate depth and surface normal predictions.
Last Context Frame Future Frame VISTA (Output)
Refer to caption Refer to caption Refer to caption
Ours Oracle VISTA
Segm. Refer to caption Refer to caption Refer to caption
Depth Refer to caption Refer to caption Refer to caption
Surf. Normals Refer to caption Refer to caption Refer to caption
(a) Short-Term
Last Context Frame Future Frame VISTA (Output)
Refer to caption Refer to caption Refer to caption
Ours Oracle VISTA
Segm. Refer to caption Refer to caption Refer to caption
Depth Refer to caption Refer to caption Refer to caption
Surf. Normals Refer to caption Refer to caption Refer to caption
(b) Mid-Term
Figure 5: Visualization of future predictions for semantic segmentation, depth, and surface normals. The illustrated scene is Frankfurt (01 (017082-017111)). Both DINO-Foresight and Vista capture the static elements effectively when the vehicle is moving forward, highlighting their capabilities in such scenarios. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Last Context Frame Future Frame VISTA (Output)
Refer to caption Refer to caption Refer to caption
Ours Oracle VISTA
Segm. Refer to caption Refer to caption Refer to caption
Depth Refer to caption Refer to caption Refer to caption
Surf. Normals Refer to caption Refer to caption Refer to caption
(a) Short-Term
Last Context Frame Future Frame VISTA (Output)
Refer to caption Refer to caption Refer to caption
Ours Oracle VISTA
Segm. Refer to caption Refer to caption Refer to caption
Depth Refer to caption Refer to caption Refer to caption
Surf. Normals Refer to caption Refer to caption Refer to caption
(b) Mid-Term
Figure 6: Visualization of future predictions for semantic segmentation, depth, and surface normals. The illustrated scene is Munster (15). Vista struggles with moving objects and smaller structures, whereas DINO-Foresight maintains more detailed predictions. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Context Frames
$X_{t-9}$ $X_{t-6}$ $X_{t-3}$ $X_{t}$
Refer to caption Refer to caption Refer to caption Refer to caption
Predicted Frames
$X_{t+3}(0.18s)$ $X_{t+6}(0.36s)$ $X_{t+9}(0.54s)$ $X_{t+12}(0.72s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
$X_{t+15}(0.9s)$ $X_{t+18}(1.08s)$ $X_{t+21}(1.26s)$ $X_{t+24}(1.44s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
$X_{t+27}(1.62s)$ $X_{t+30}(1.8s)$ $X_{t+33}(1.98s)$ $X_{t+36}(2.16s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
$X_{t+39}(2.34s)$ $X_{t+42}(2.52s)$ $X_{t+45}(2.7s)$ $X_{t+48}(2.88s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Long-term semantic segmentation, depth and surface normal predictions. The illustrated scene is Frankfurt (01 (011791-011820)). DINO-Foresight consistently preserves motion dynamics and intricate details in complex scenes over extended time horizons. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Context Frames
$X_{t-9}$ $X_{t-6}$ $X_{t-3}$ $X_{t}$
Refer to caption Refer to caption Refer to caption Refer to caption
Predicted Frames
$X_{t+3}(0.18s)$ $X_{t+6}(0.36s)$ $X_{t+9}(0.54s)$ $X_{t+12}(0.72s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
$X_{t+15}(0.9s)$ $X_{t+18}(1.08s)$ $X_{t+21}(1.26s)$ $X_{t+24}(1.44s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
$X_{t+27}(1.62s)$ $X_{t+30}(1.8s)$ $X_{t+33}(1.98s)$ $X_{t+36}(2.16s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
$X_{t+39}(2.34s)$ $X_{t+42}(2.52s)$ $X_{t+45}(2.7s)$ $X_{t+48}(2.88s)$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Long-term semantic segmentation, depth and surface normal predictions. The illustrated scene is Frankfurt (01 (006570-006599)). DINO-Foresight excels in predicting the motion of the nearby car but faces challenges with distant, low-motion objects, highlighting areas for future improvement. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Appendix B Limitations and Future Work
In this work, we introduced DINO-Foresight, a simple yet effective method for semantic future prediction based on forecasting VFM features. While the results are promising, several exciting research opportunities remain.
First, our approach uses a simple masked transformer architecture with a SmoothL1 loss, which is deterministic and does not account for the inherent uncertainty in future predictions. Incorporating stochastic elements could better capture this ambiguity, leading to more robust predictions. Although we explored strategies to reduce training compute for high-resolution feature prediction, inference-time compute demands remain unchanged. Future work could address this by adopting hierarchical transformer architectures [72], which would not only improve efficiency but also enable the model to handle even higher feature resolutions.
To reduce the dimensionality of multi-layer VFM features, we relied on PCA, a linear projection method. While effective, exploring advanced dimensionality reduction techniques, such as VAEs, could further enhance feature representation and boost performance.
Another promising direction lies in leveraging the simple and scalable design of DINO-Foresight to investigate its behavior with larger datasets and bigger models. Our strong results on Cityscapes suggest that scaling both the data and model size, including the VFM encoder and prediction transformer, could unlock further improvements in future forecasting.
Finally, extending DINO-Foresight to incorporate action conditioning and action prediction presents an exciting opportunity to transform it into a world model capable of reasoning and control.
Overall, these research directions highlight the flexibility and growth potential of our approach, paving the way for further advancements in semantic future prediction.
Appendix C Implementation Details of Prediction Heads
We provide implementation details for the heads trained on different downstream tasks. The DPT head is used for semantic segmentation, depth estimation, and surface normal estimation. We adopt the DPT [70] implementation from Depth Anything [93, 94], setting the feature dimensionality to 256 and configuring dptoutchannels = [128, 256, 512, 512]. For all tasks, models are trained for 100 epochs with a batch size of 128 ($16\times 8$ GPUs). The learning rate is set to 0.0016, using the AdamW optimizer with linear warmup for the first 10 epochs, and weight decay is 0.0001. For semantic segmentation, we use a polynomial scheduler and cross-entropy loss with 19 classes. For depth estimation, we use a cosine annealing scheduler and cross-entropy loss, with 256 classes. For surface normal estimation, we employ a polynomial scheduler and a loss function combining cosine similarity and $L_{2}$ loss with weighted averaging, using 3 classes.
For the Mask2Former head used in instance segmentation, we implement our approach using the official Mask2Former [17] and Detectron2 [90] codebases. The main difference compared to the official Mask2Former configuration for Cityscapes instance segmentation is the input feature maps. In our approach, the four multi-scale feature maps expected by Mask2Former are derived from the PCA features. These PCA features are first projected to 128, 256, 512, and 1024 dimensions and then resized so their spatial resolutions are ×$\times 4$, ×$\times 2$, ×$\times 1$, and ×$\times 0.5$ relative to the original resolution of the DINOv2 ViT-B outputs. We train using the AdamW optimizer, with a batch size of 64 ($8\times 8$ GPUs), learning rate of 0.00032, weight decay of 0.05, and 67,500 iterations, with a polynomial scheduler.